OpenStudy (anonymous):

Zach ran t times around the track at school, which is 1 fourth of MILE mile. He ran a total of 9 TENTHS mile. Solve for t.

OpenStudy (anonymous):

Well since zach ran at least .75 we know he went 3 laps already. but he ran .9 miles so .9-.75=.15miles since each lap is .25 miles he ran 3+(.15/.25) miles = 3+0.6 = 3.6 miles

OpenStudy (anonymous):

in other words t = total distance ran divided by 1 quarter of a mile

OpenStudy (anonymous):

So simply put t=\[\frac{ .9 }{ .25 }\] which =3.6 miles